answersLogoWhite

0

You can get 8 with (2, 6), (3, 5), (4, 4), (5, 3) or (6, 2), each with a probability of (1/6)*(1/6) = 1/36

The the total probability is 5/36 = 0.1388... recurring.

User Avatar

Wiki User

∙ 12y ago
User Avatar

What else can I help you with?

Related Questions

When two dice are rolled what is the probability of getting a sum of 14?

You can't get 14 with two regular six-sided dice ! The highest you can get with one throw is 12.


If 11 is the sum and you roll two dice what is the probability of getting 11 if you toss the dice 54 times?

The probability of getting 11 with one throw of 2 dice is 1/6*1/6*2 = 1/18 So the probability of not getting 11 with 1 throw of the dice is 17/18. Tossing the dice 54 times, the probability of not getting 11 54 times is (17/18)54 = 0.0456... So the probability of at least 1 roll of 11 is 1 - 0.0456 = 0.954


What is the probability that the sum is not seven if you throw 2 dice?

It is 30/36 = 5/6


What are the chances of getting a sum of 70 on two dice?

The probability is zero. The largest possible sum on two dice is 12.


When two dice are rolled find the probability of getting a sum of greater than 15?

Two standard dice cannot have a sum greater than 12. The probability, then, of getting a sum greater than 15 is zero.


When two dice are rolled find the probability of getting a sum less than 13?

Assuming these are regular dice, the probability is 1.


What is the probability of obtaining a sum of 11 in a single throw with 3 dice?

The probability is 27/216 = 3/8 = 0.125


What is the probability of getting the sum of 8 from a pair of dice?

It is 5/36.


When two dice are rolled what is the probability of not getting a sum of 12?

The probability of rolling 12 with 2 dice is 1 in 36. The probability of not rolling 12 with 2 dice is 35 in 36.


When two dice are rolled the probability of getting a sum of 16?

With two normal six-sided dice, the maximum sum can only be 12 (6 x 2 dice), so there is zero probability of rolling a sum that adds to 16.


How many times must you throw two standard dice to be sure you get the same sum at least twice?

The probability of gettting a particular sum on a standard set of dice depends on which sum you are seeking. For example, the sums of 2 and 12 have a probability of 1 in 36, or about 0.0278; while the sum of 7 has a probability of 6 in 36, or 1 in 6, or about 0.167.Specifically answering the question; it is not possible to guarantee a particular outcome in a random throw, or in a series of random throws, of the dice. You can only talk about probability. Let's take the worst case of trying to throw a 2 or a 12. Even of you throw the dice 100 times, the probability is only 0.0278100, or about 2.34 x 10-156 that you will not throw the 2 or 12; so, the probability is extremely good that you will throw a 2 or 12 in 100 throws, but it is not guaranteed. That's the thing about probability.


When two dice are rolled find the probability of getting a 6 or 7?

Probability that the sum is 6 = 5/36 Probability that the sum is 7 = 6/36

Trending Questions
How many ways can you arrange 5 books on a shelf order is important? What is the cause of the peak in the graph line at 12 pm? What kind diagram is used to graph two-variable data for correlation analysis? What are The University of Hawaii-West Oahu's 25th-75th percentile ACT scores? There were 6 purple socks and 4 orange socks in a drawer Zucky picked one sock without looking and then another without looking replacing the first what is the probability that he picked 2 purple? If a bag contains 4 green marbles an 4 yellow marbles 5 white marbles what would be more likely to pull a red marble or a white marble from the bag? The process of using sample statistics to draw conclusions about true population parameters? What percent of iowans vote? What is the average size of a megabat? What is the drop rate of the inquisitor guard armor in AQWorlds? What is the median in a set of numbers? What does it mean to be a doge? What is the 4.5 percent return for 500000? What is the average number of trades a day trader executes in a day? What shows how the data set is spread around the mean? What is a pairing of each number in a given set of numbers with exactly one number in another set? Why did the women have a better chance of surviving the titanic than men? When there are no identifiable trends in a graph it means there is no relationship between the two variables? If you had two periods the month before and the next month your stomach gets hard and you have synthems of pregnancy what does that mean? What does bunglow mean?